OpenStudy (anonymous):

What is the sum of the geometric sequence 1, –4, 16, … if there are 7 terms?

OpenStudy (anonymous):

do you know the formula?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can you write it please? :)

OpenStudy (zepp):

You can write a geometric sequence like this: {a, ar, ar^2, ar^3, ar^4...}

OpenStudy (zepp):

All you need to find is the common ratio, which is -4 Explanation: -4/1 = -4 16/-4 = -4 r would be -4 The first term is 1 So your sequence should look like this {1, 1*-4, 1*(-4)^2....}

OpenStudy (zepp):

To get the sum, use \[\sum_{a=1}^{7}ar^k\]

OpenStudy (anonymous):

\[\LARGE r={a_n\over a_{n-1}}={a_2\over a_1}={-4\over 1}=-4\] now substitute: \[\LARGE S_7=a_1\cdot \frac{1-r^{7}}{1-r}\]

OpenStudy (zepp):

How do you make the text bigger? @Kreshnik D:

OpenStudy (anonymous):

\[\Huge \text{I DON'T KNOW!}\] \[\Huge 1+2=3\] :P use \LARGE x-2 ... or \huge 45654 or the one I just used:" \Huge x-26565445698765435357357357357... :)

OpenStudy (zepp):

Alright thanks :D \[\huge \sum_{a=1}^{7}ar^k = a(\frac{1-r^7}{1-r})=1(\frac{1-(-4)^7}{1-(-4)})\] Cool ^^

OpenStudy (anonymous):

@zepp which one's better? \[ \Huge \left( \frac32 \right) \] \[\Huge (\frac32) \] ?? use: \left( \frac32 \right)

OpenStudy (zepp):

First one :D

OpenStudy (anonymous):

\[\Huge \lim_{x\to \infty}\left[\frac{x^2+2x}{\sqrt{x^7-56x}}\right]\] this looks cute :) lol

OpenStudy (zepp):

Indeed ;D @Kreshnik